OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 7x + 8 = 0. Estimate irrational solutions to the nearest tenth. A. {–12, 5} B. {–11.1, –2.9} C. {–5.6, –1.4} D. {–3.9, 12.1}

OpenStudy (michele_laino):

hint: \[\Large x = \frac{{ - 7 \pm \sqrt {{7^2} - 4 \times 1 \times 8} }}{{2 \times 7}}\]

OpenStudy (anonymous):

For quadratic formulas you have the basic form \[ax^2+bx+c=0\] And than you should determine your determinator which is \[b^2-4ac\] And then your solutions are \[\frac{ -b \pm \sqrt{determinator} }{ 2a }\]

OpenStudy (anonymous):

i need help on the solving 

OpenStudy (anonymous):

i already had help at the begenning but, the user left off at the following part

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

that's right! So we have: \[\begin{gathered} {x_1} = \frac{{ - 7 + 4.12}}{{14}} = - 0.21 \hfill \\ \hfill \\ {x_2} = \frac{{ - 7 - 4.12}}{{14}} = - 0.79 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. I have made an error, here are the right formulas:

OpenStudy (michele_laino):

\[\Large \begin{gathered} x = \frac{{ - 7 \pm \sqrt {{7^2} - 4 \times 1 \times 8} }}{2} \hfill \\ {x_1} = \frac{{ - 7 + 4.12}}{2} = - 1.44 \hfill \\ \hfill \\ {x_2} = \frac{{ - 7 - 4.12}}{2} = - 5.56 \hfill \\ \end{gathered} \]